1.Using the data in Table 1, calculate the percentage of Low Birth Weight for African American and White infants in Minnesota during 2015.

Table 1. Data for African American and White Births, Minnesota 2015 All Births (number) 49,726 Non-Hispanic White Non-Hispanic African American/ 7,966 Black MN Total LBW 750 4,503 2,881 All Births in Minnesota 2015-not limited to singleton births. Data source: Minnesota Department of Health, Community and Family Health Division
Answer #1

Percentage of low birth weight = ( number of low birth weight / number of births ) × 100
Percentage of low birth weight for white are = ( 2881/49726) × 100 = 5.79 %
Percentage of low birth weight for African American = (750/7966) × 100 = 9.41 %